OpenStudy (anonymous):

The owner of a bike shop sells unicycles and bicycles and keeps inventory by counting seats and wheels. One day, she counts 20 seats and 28 wheels. Which equation in slope-intercept form shows that the number of bicycles (x) plus the number of unicycles (y) is 20? A. y = -x - 20 B. y = x - 20 C. y = x + 20 D. y = -x + 20

OpenStudy (daniellelovee):

use the same method you used to answer the other question

OpenStudy (daniellelovee):

x+y=20 x+2y=28

OpenStudy (anonymous):

would it be C?

OpenStudy (daniellelovee):

hold on

OpenStudy (daniellelovee):

i think so because the other question was asking for the wheels but this is not specifying soo

OpenStudy (daniellelovee):

because if we work x+y=20 backwards we get 20-x=y

OpenStudy (daniellelovee):

which will give -x+20=y makes sense to you?

OpenStudy (anonymous):

i think so.

OpenStudy (daniellelovee):

lol did you understand what I said?? if you didn't tell me where you got lost

OpenStudy (daniellelovee):

in my opinion the answer is D

OpenStudy (anonymous):

You think its D?

OpenStudy (anonymous):

Shall II go with D?

OpenStudy (daniellelovee):

your choice lol

OpenStudy (daniellelovee):

i told you why i think is D

OpenStudy (anonymous):

Correct you are fellow Jeti warrior. xD LOL

OpenStudy (daniellelovee):

I was right? :D

OpenStudy (daniellelovee):

thats great lol because I was doubting myself

OpenStudy (anonymous):

yes